two infinitely long uniformly charged ( having linear charge densities t1 and t2 respectively ) parallel straight wires are kept r distance apart. what is the force per unit length on that wire?

Grade:11

1 Answers

Abhishek Singh
7 years ago
let consider a segment on second wire of length l
so the field caused by the other one  is\frac{t1}{4\pi\varepsilon r }
so the for force  = charge * field = field * t2*l
force per unit lengt = field *t2 = \frac{t1*t2}{4\pi\varepsilon r }